# Robust Equilibria in Shared Resource Allocation via Strengthening Border’s Theorem

### v1

@@ Abstract
We consider repeated allocation of a shared resource via a non-monetary mechanism, wherein a single item must be allocated to one of multiple agents in each round. We assume that each agent has i.i.d. values for the item across rounds, and additive utilities. Past work on this problem has proposed mechanisms where agents can get one of two kinds of guarantees: $(i)$ (approximate) Bayes-Nash equilibria via linkage-based mechanisms which need extensive knowledge of the value distributions, and $(ii)$ simple distribution-agnostic mechanisms with robust utility guarantees for each individual agent, which are worse than the Nash outcome, but hold irrespective of how others behave (including possibly collusive behavior). Recent work has hinted at barriers to achieving both simultaneously. Our work however establishes this is not the case, by proposing the first mechanism in which each agent has a natural strategy that is both a Bayes-Nash equilibrium and also comes with strong robust guarantees for individual agent utilities.
Our mechanism comes out of a surprising connection between the online shared resource allocation problem and implementation theory. In particular, we show that establishing robust equilibria in this setting reduces to showing that a particular subset of the Border polytope is non-empty. We establish this via a novel joint Schur-convexity argument. This strengthening of Border’s criterion for obtaining a stronger conclusion is of independent technical interest, as it may prove useful in other settings.
@@ 1 Introduction
Consider a single indivisible public resource being repeatedly allocated between multiple agents – for example, a scientific instrument shared by multiple university labs. Since the resource is public, its allocation should be determined ideally without using monetary transfers. Moreover, the principal wants to allocate the resource in a way that is both efficient (so labs get the instrument only when they have great need for it) and also fair (so that the amount of time each lab gets to use it is roughly proportional to some pre-determined share). Each individual agent is of course self-serving, and so some mechanism is required to encourage agents to request for the resource only when they need it the most. And in the absence of money, the principal may not be able to verify agents’ reports, or have any strong beliefs about their valuations and actions.
The above problem was first studied under a simple model by [guo2009competitive], with a single indivisible item per round, and agents with random valuations across rounds. Early work on this setting [guo2009competitive, cavallo2014incentive, balseiro2019multiagent, gorokh2021monetary] focused on the question of how efficient a non-monetary mechanism could be in such a setting; as we discuss in Section 1.3, this led to several mechanisms with near-efficient Bayes-Nash equilibria. All these mechanisms, however, critically depend on knowing the exact value distributions.
[…]
So is it possible to design mechanisms that unite the above two streams for the setting of [guo2009competitive]? We define this formally in Section 2.2, but at a high level, we seek mechanisms that support a good robust equilibrium: a strategy that simultaneously satisfies the following:
@@
This definition is of course under-specified in terms of what we mean by ‘good’: note that never allocating, or allocating uniformly at random, supports any strategy both as an equilibrium, and also gives every agent a (weak) robust guarantee. Our main result however is a new mechanism for repeated non-monetary allocation which admits a simple strategy that is simultaneously a Bayes-Nash equilibrium in the infinite-horizon limit (and close to Bayes-Nash equilibrium in the finite case), and also matches the current best robustness guarantee. Moreover, we achieve this via a surprising strengthening of Border’s theorem, which may prove useful in other settings.
@@ 1.1 Overview of our Mechanism and Main Result
[…]
@@
Within this backdrop, we propose a new mechanism we call Budgeted Robust Border (BRB) (see Mechanism 1). At a high level, the mechanism has two simple components:
@@
For the exact expression for the budget $B_{i}$ see Proposition 4.1; the additional $o(T)$ budget is needed to get high probability guarantees. The allocation probabilities $\{p_{i}^{S}\}$ are more complex (and indeed, that is where the main novelty of our work lies, as we discuss next). However, at this point, we can already summarize our main result, as follows (see Theorem 4.3):
Under the BRB Mechanism, the policy where player $i$ bids in round $t$ whenever her value $V_{i}[t]$ is in the top $\alpha_{i}$-quantile (and subject to her budget), is simultaneously: Robust with factor $\lambda_{\mathtt{ROB}}\geq 1/2$ A Bayes-Nash equilibrium with $\lambda_{\mathtt{NASH}}=1-\prod_{j=1}^{n}(1-\alpha_{j})\geq 1-1/e$
In Lemma 2.1 we prove the above $\lambda_{\mathtt{NASH}}$ is minimax optimal for any vector of fair shares $\{\alpha_{i}\}_{[n]}$; $\lambda_{\mathtt{ROB}}=1/2$ matches the current best [gorokh2019remarkable, banerjee2023robust, fikioris2023online].
@@ 1.2 Overview of our Techniques
Our main conceptual idea is a connection between repeated non-monetary allocation and implementation theory (in particular, Border’s theorem). Our main technical novelty is a way to modify the Border flow network in order to select $1/2$-robust equilibria. Our mechanism uses three main ideas: $(i)$ a repeated all-pay mechanism to control bidding rates, $(ii)$ guaranteeing good interim allocations via the regular Border condition, and $(iii)$ strengthening the Border condition to guarantee robustness. We now describe these in brief.
@@
We start by first trying to get good equilibria. In the infinite horizon setting, our idea is as follows: We restrict agents to only bid or not in each round, and further restrict them to bid at most an $\alpha_{i}$ fraction of the time. Our mechanism aims to guarantee agent $i$ a good interim allocation conditioned on them bidding (the probability that they are allocated), assuming every other agent $j$ bids independently with probability at most $\alpha_{j}$. Finally, to pass to the finite horizon, we use an idea from [gorokh2021monetary] and employ an all-pay mechanism with budgets $\{\alpha_{i}T+o(T)\}$, to ensure that agents are not incentivized to bid at a rate higher than $\alpha_{i}$, while also not running out of budget with high probability.
@@
To ensure good interim allocations, we use probabilistic allocation with preset probabilities $p_{i}^{S}$ for assigning the item to agent $i$ when the set $S$ of agents bid for each subset $S\subseteq[n]$, and agents $i\in S$. These are chosen to ensure each agent’s bids are accepted with probability at least $\lambda_{\mathtt{NASH}}$, assuming each agent $i$ bids independently with probability $\alpha_{i}$. The bid budgets and probabilistic allocation serves to somewhat insulate each agent from the actions of others. Consequently, in equilibrium, agent $i$’s best response is to bid when her value is in the highest $\alpha_{i}$ quantile, and which makes bids independent across agents (since valuations are independent). Border’s theorem (see Theorem 5.2) provides necessary and sufficient conditions for which interim allocation probabilities are feasible; in particular, we use this to show we can always guarantee $\lambda_{\mathtt{NASH}}\geq 1-\prod_{j}(1-\alpha_{j})\geq(1-1/e)\approx 0.63.$ (see Theorems 4.3 and 4.2). Our guarantee is in fact minimax optimal for any vector of fair shares $\{\alpha_{i}\}_{[n]}$: in Lemma 2.1 we show a valuation profile that gives a matching upper bound even if the principal knows agents’ realized values.
@@
The challenge however is that many interim allocation probabilities admitted by Border’s theorem are not robust: if all other agents collude against agent $i$, they can severely limit agent $i$’s allocation. In fact, different allocation rules inducing the same interim allocation can have vastly different robustness guarantees (see Section 6.2). Even computing $\lambda_{\mathtt{ROB}}$ for a given allocation rule appears difficult, as we need to search over all potentially correlated bidding schemes by the other agents.
@@
Our main technical novelty is in identifying the subset of interim allocations allowed by Border’s theorem which are $1/2$-robust. To do this, we consider any chosen agent $i$, and characterize the bang-per-buck (in terms of blocking agent $i$ per bid token spent) of each possible set $S\subseteq[n]\setminus\{i\}$ of colluding agents bidding in a round. Using this, we show that the critical case is allocating when only 2 agents bid (when agents collude, they can make this the dominant case). We show that it is possible to make the mechanism $1/2$-robust (matching the current best robustness guarantee), by modifying the Border flow network by adding appropriate bounds to edges going out of each doubleton subset $\{i,j\}\subset[n]$ (Fig. 2). This leads to a modification of Border’s criterion for ensuring the resultant network still supports the same maximum amount of flow (Lemma 6.2). Finally, in Lemma 6.3, we demonstrate the resultant polytope is non-empty, via establishing the Schur-convexity/concavity of a certain bivariate function that comes out of our new criterion.
@@
Taken together, our arguments demonstrate allocation probabilities that simultaneously guarantee for all agents at least an $(1-1/e)$ fraction of their ideal utility at Nash equilibrium and, at the same time, a $1/2$ fraction of their ideal utility even if other agents collude. While the allocation rules in the general case are somewhat involved, we illustrate them explicitly in two simple corner cases – $2$ agents with arbitrary shares in Section 3, and $n$ agents with equal shares in Section 6.
@@ 1.3 Related Work
[…]
@@ 2 Preliminaries
2.1 Model
We consider the following simple setting of repeatedly allocating a single, indivisible item among $n$ agents; this was first introduced in the work of [guo2009competitive], and has since been widely studied, as we discuss in Section 1.3. At each discrete time-step $t=1,2,\dots,T$, a principal receives a new item and must select an agent $i\in[n]$ to allocate the item to (or not allocate the item).
@@ 2.2 Ideal Utility and Benchmarks
As in previous work in this setting [gorokh2019remarkable, banerjee2023robust, fikioris2023online], we assume each agent $i$ has an exogenously defined fair share $\alpha_{i}$, where each $\alpha_{i}$ is nonnegative and $\sum_{i=1}^{n}\alpha_{i}=1$. The fair share $\alpha_{i}$ of agent $i$ represents the fraction of rounds we want agent $i$ to win the item under ideal circumstances. As done by previous work, we use the benchmark of ideal utility. The ideal utility of an agent $i$ is the maximum expected utility they can obtain from a single round if they can obtain the item simply by requesting it, but they are only allowed to request it with probability at most their fair share $\alpha_{i}$. Formally, the (per-round) ideal utility is the following.
@@
[…]
We study what fraction of her ideal utility an agent $i$ can guarantee in two settings: First, we examine the performance of Bayes-Nash equilibria, where each agent is trying to maximize their own utility. Second, we examine what fraction of her ideal utility agent $i$ can guarantee robustly, i.e., even if the other agents $j\neq i$ are playing adversarially and collude to harm agent $i$ without regard for their own utilities. Critically, we want an agent to achieve both the above with the same (ideally simple) strategy. As in Section 1, we want a mechanism that achieves the following.
$\lambda_{\mathtt{NASH}}$-Nash $\lambda_{\mathtt{ROB}}$-robust equilibria: A mechanism with policy profile $(\pi_{1},\pi_{2},\ldots,\pi_{n})$ s.t. $(\pi_{1},\pi_{2},\ldots,\pi_{n})$ forms a Nash equilibrium as $T\to\infty$ in which each agent $i$ gets a $\lambda_{\mathtt{NASH}}$ fraction of their ideal utility. Policy $\pi_{i}$ is $\lambda_{\mathtt{ROB}}$-robust with $\lambda_{\mathtt{ROB}}$ is not much smaller than $\lambda_{\mathtt{NASH}}$: it guarantees agent $i$ a $\lambda_{\mathtt{ROB}}$ fraction of her ideal utility even if agents $j\neq i$ collude and act adversarially.
To understand how the ideal utility benchmark behaves under Bayes-Nash equilibria, we first present a minimax bound on what fraction $\lambda_{\mathtt{NASH}}$ of ideal utility agents can get under ideal conditions. This will help later certify the minimax optimality of our mechanism in the equilibrium setting. Note that this bound is mechanism-agnostic, as in our example below, the mechanism has access to the agents’ realized values.

### v2

@@ Abstract
We consider repeated allocation of a shared resource via a non-monetary mechanism, wherein a single item must be allocated to one of multiple agents in each round. We assume that each agent has i.i.d. values for the item across rounds, and additive utilities. Past work on this problem has proposed mechanisms where agents can get one of two kinds of guarantees: $(i)$ (approximate) Bayes-Nash equilibria via linkage-based mechanisms which need extensive knowledge of the value distributions, and $(ii)$ simple distribution-agnostic mechanisms with robust utility guarantees for each individual agent, which are worse than the Nash outcome, but hold irrespective of how others behave (including possibly collusive behavior). Recent work has hinted at barriers to achieving both simultaneously. Our work however establishes this is not the case, by proposing the first mechanism in which each agent has a natural strategy that is both a Bayes-Nash equilibrium and also comes with strong robust guarantees for individual agent utilities. Our mechanism comes out of a surprising connection between the online shared resource allocation problem and implementation theory, and uses a surprising strengthening of Border’s theorem. In particular, we show that establishing robust equilibria in this setting reduces to showing that a particular subset of the Border polytope is non-empty. We establish this via a novel joint Schur-convexity argument. This strengthening of Border’s criterion for obtaining a stronger conclusion is of independent technical interest, as it may prove useful in other settings.
@@ 1 Introduction
Consider a single indivisible public resource being repeatedly allocated between multiple agents – for example, a scientific instrument shared by multiple university labs. Since the resource is public, its allocation should be determined ideally without using monetary transfers. Moreover, the principal wants to allocate the resource in a way that is both efficient (so labs get the instrument only when their need is great) and fair (so that the amount of time each lab gets to is roughly proportional to some pre-determined share). Each individual agent is of course self-serving, and so some mechanism is required to encourage agents to request for the resource only when they need it the most. And in the absence of money, the principal may not be able to verify agents’ reports, or have any strong beliefs about their valuations and actions.
The above problem was first studied under a simple model by [guo2009competitive], with a single indivisible item per round, and agents with random valuations across rounds. Early work on this setting [guo2009competitive, cavallo2014incentive, balseiro2019multiagent, gorokh2021monetary] focused on the question of how efficient a non-monetary mechanism could be in such a setting; as we discuss in subsection˜1.3, this led to several mechanisms with near-efficient Bayes-Nash equilibria. All these mechanisms, however, critically depend on knowing the exact value distributions.
More recently, [gorokh2019remarkable] pointed out that in non-monetary settings, the inability to make interpersonal comparisons makes the knowledge of exact value distributions unlikely. This led to a line of work [gorokh2019remarkable, banerjee2023robust, fikioris2023online] focusing on mechanisms that, under minimal assumptions, have robust individual-level guarantees, which hold irrespective of how other agents behave. On the positive side, such guarantees adhere closer to Wilson’s doctrine that mechanisms should be as ‘detail-free’ (i.e., distribution agnostic) as possible [wilson1985game]. On the other hand, it is unclear how predictive such results are of agent behavior; these works do not show if their mechanisms admit any simple equilibria, and in general, this is a difficult task in any repeated mechanism. More surprisingly, it was recently reported [chido] that in one of the simplest such mechanisms called Dynamic Max-Min Fair sharing (DMMF), there is in fact no equilibrium under a natural class of strategies (roughly speaking, where each agent’s actions only depend on their value in each round). While this result involves a technical characterization of a certain high-dimensional Markov chain induced under DMMF, it appears to support a natural critique of robustness results: that they involve strategies that are too pessimistic, and hence not supported in an equilibrium.
So is it possible to design mechanisms that unite the above two streams for the setting of [guo2009competitive]? We define this formally in subsection˜2.1, but at a high level, we seek mechanisms that support a good robust equilibrium: a strategy that simultaneously satisfies the following:
@@
This definition is of course under-specified in terms of what we mean by ‘good’: note that never allocating, or allocating uniformly at random, supports any strategy both as an equilibrium, and also gives every agent a (weak) robust guarantee. Our main result however is a new mechanism for repeated non-monetary allocation which admits a simple strategy that is a Bayes-Nash equilibrium in the infinite-horizon limit (and close to Bayes-Nash equilibrium in the finite case) with high equilibrium and robust performance. Moreover, we achieve this via a surprising strengthening of Border’s theorem, which may prove useful in other settings.
@@ 1.1 Overview of our Mechanism and Main Result
We consider the setting of [guo2009competitive], with a horizon of $T$ rounds, a single indivisible item to allocate per round, and $n$ agents with random valuations. Agent $i$’s values $V_{i}[t]$ are i.i.d. across rounds and independent across agents. Building on the ideas of [gorokh2019remarkable], we eschew efficiency to focus on share-based guarantees: each agent $i$ has a pre-determined fair share $\alpha_{i}$ (with $\sum_{i}\alpha_{i}=1$), which then allows us to define a per-user utility benchmark, rather than compare utilities across users. With an $\alpha_{i}$ share, agent $i$’s best hope is to get her favorite $\alpha_{i}T$ rounds (i.e., rounds $t\in[T]$ where $V_{i}[t]$ is in the top $\alpha_{i}$ quantile of her value distribution). Formally (Definition˜2.1), an agent’s ideal utility $v_{i}^{\star}$ with fair share $\alpha_{i}$ is the maximum utility she can get if awarded the resource with probability at most $\alpha_{i}$.
@@
Within this backdrop, we propose a new mechanism we call Budgeted Robust Border (BRB) (see Mechanism˜1). At a high level, the mechanism has two simple components:
@@
For the exact expression for the budget $B_{i}$ see Proposition˜3.3; the additional $o(T)$ budget is needed to get high probability guarantees. The allocation probabilities $\{p_{i}^{S}\}$ are more complex (and indeed, that is where the main novelty of our work lies, as we discuss next). However, at this point, we can already summarize our main result, as follows (see Theorem˜3.6):
Under the BRB Mechanism, the policy where player $i$ bids in round $t$ whenever her value $V_{i}[t]$ is in the top $\alpha_{i}$-quantile (and subject to her budget), is simultaneously: Robust with factor $\lambda_{\mathtt{ROB}}\geq 1/2$ A Bayes-Nash equilibrium with $\lambda_{\mathtt{NASH}}=1-\prod_{j=1}^{n}(1-\alpha_{j})\geq 1-1/e\approx 0.63$
In Lemma˜2.2, we prove the above $\lambda_{\mathtt{NASH}}$ is minimax optimal for any vector of fair shares $\{\alpha_{i}\}_{[n]}$. $\lambda_{\mathtt{ROB}}=1/2$ matches the one in [gorokh2019remarkable, banerjee2023robust, fikioris2023online], which was the best-known until quite recently, when [lin2025online] provided a more complicated strategy in the mechanism of [gorokh2019remarkable] that obtains $\lambda_{\mathtt{ROB}}=2-\sqrt{2}\approx 0.59$. However, everyone playing their robust strategy is not an equilibrium. Moreover, their mechanism is more complicated, so equilibrium behavior is hard to analyze, and only robustness guarantees are known.
@@ 1.2 Overview of our Techniques
Our main conceptual idea is a connection between repeated non-monetary allocation and implementation theory (particularly, Border’s theorem). Our main technical novelty is modifying the Border flow network to select $1/2$-robust equilibria. This strengthening of Border’s theorem to show that this restricted flow polytope is also guaranteed to be non-empty can be of independent interest. Our mechanism uses three main ideas: $(i)$ a repeated all-pay mechanism to control bidding rates, $(ii)$ guaranteeing good interim allocations via the regular Border condition, and $(iii)$ strengthening the Border condition to guarantee robustness. We now briefly describe these.
@@
We start by first trying to get good equilibria. In the infinite-horizon setting, our idea is as follows: We restrict agents to only bid or not in each round, and further restrict them to bid at most an $\alpha_{i}$ fraction of the time. Our mechanism aims to guarantee agent $i$ a good interim allocation conditioned on them bidding (the probability that they are allocated), assuming every other agent $j$ bids independently with probability at most $\alpha_{j}$. Finally, to pass to the finite horizon, we use an idea from [gorokh2021monetary] and employ an all-pay mechanism with budgets $\{\alpha_{i}T+o(T)\}$, to ensure that agents are not incentivized to bid at a rate higher than $\alpha_{i}$, while also not running out of budget with high probability.
@@
To ensure good interim allocations, we use probabilistic allocation with preset probabilities $p_{i}^{S}$ for assigning the item to agent $i$ when the set $S$ of agents bid for each subset $S\subseteq[n]$, and agents $i\in S$. These are chosen to ensure each agent’s bids are accepted with probability at least $\lambda_{\mathtt{NASH}}$, assuming each agent $i$ bids independently with probability $\alpha_{i}$. The bid budgets and probabilistic allocation serve to somewhat insulate each agent from the actions of others. Consequently, in equilibrium, agent $i$’s best response is to bid when her value is in the highest $\alpha_{i}$ quantile, and which makes bids independent across agents (since valuations are independent). Border’s theorem (see Theorem˜4.2) provides necessary and sufficient conditions for which interim allocation probabilities are feasible; in particular, we use this to show we can always guarantee $\lambda_{\mathtt{NASH}}\geq 1-\prod_{j}(1-\alpha_{j})\geq(1-1/e)\approx 0.63.$ (see Theorems˜3.6 and 3.5). Our guarantee is in fact minimax optimal for any vector of fair shares $\{\alpha_{i}\}_{[n]}$: in Lemma˜2.2 we show a valuation profile that gives a matching upper bound even if the principal knows agents’ realized values.
@@
The challenge however is that many interim allocation probabilities admitted by Border’s theorem are not robust: if all other agents collude against agent $i$, they can severely limit agent $i$’s allocation. In fact, different allocation rules inducing the same interim allocation can have vastly different robustness guarantees (see subsection˜5.2). Even computing $\lambda_{\mathtt{ROB}}$ for a given allocation rule appears difficult, as we need to search over all potentially correlated bidding schemes by the other agents.
@@
Our main technical novelty is in identifying the subset of interim allocations allowed by Border’s theorem which are $1/2$-robust. To do this, we consider any chosen agent $i$, and characterize the bang-per-buck (in terms of blocking agent $i$ per bid token spent) of each possible set $S\subseteq[n]\setminus\{i\}$ of colluding agents bidding in a round. Using this, we show that the critical case is allocating when only 2 agents bid (colluding agents can always make this the dominant case). We show that it is possible to make the mechanism $1/2$-robust (matching [gorokh2019remarkable, banerjee2023robust, fikioris2023online]), by modifying the Border flow network by adding appropriate bounds to edges going out of each doubleton subset $\{i,j\}\subset[n]$ (Figure˜5.1). This leads to a modification of Border’s criterion for ensuring the resultant network still supports the same maximum amount of flow (Lemma˜5.2). Finally, in Lemma˜5.3, we demonstrate the resultant polytope is non-empty, via establishing the Schur-convexity/concavity of a certain bivariate function that comes out of our new criterion.
Finally, in section˜6, we demonstrate how to efficiently sample from a probabilistic allocation rule in the above polytope using the techniques of [bhalgat2013optimal]. Specifically, we design a linear program whose feasible region corresponds to the above polytope. However, a solution of this LP is of exponential size (one probabilistic allocation rule for every $S\subseteq[n]$), which means that solving it is hopeless. Instead, we calculate a dual solution of polynomial size, which we can use to calculate the desired allocation probabilities for a given set $S$. Specifically, we calculate dual solutions by running an instance of the Hedge algorithm. These dual solutions can be used to sample from allocation probabilities $p_{i}^{S}$ that result in the $1/2$-robustness and equilibrium utility guarantees that are only $O(\sqrt{\log n/K)}$ away from the optimal, where $K$ is the number of steps of the Hedge algorithm.
@@
Taken together, our arguments demonstrate allocation probabilities that simultaneously guarantee for all agents at least (11/e)(1-1/e) fraction of their ideal utility at Nash equilibrium and, at the same time, a $1/2$ fraction of their ideal utility even if other agents collude. While the allocation rules in the general case are somewhat involved, we illustrate them explicitly in two simple corner cases – $2$ agents with arbitrary shares in section˜A, and $n$ agents with equal shares in section˜5.
@@ 1.3 Related Work
Allocating shared resources without money is a foundational problem in economics. The intersection with worst-case analysis and approximation was first explored by [procaccia2009approximate], sparking renewed interest in this area. Repeated allocation of an indivisible item with random valuations was introduced by [guo2009competitive], and has since been successful in applications like course allocation [budish2017course], food banks [prendergast2022allocation], and cloud computing [vasudevan2016customizable].
Prior work in this setting falls into two main streams. The first focuses on emulating monetary mechanisms using non-monetary tools [guo2010, cavallo2014incentive, gorokh2021monetary, balseiro2019multiagent]. This culminates in the black-box reduction of [gorokh2021monetary], which uses a repeated all-pay auction to emulate any monetary mechanism with vanishing efficiency loss. Their approach builds on the idea of “linking decisions” from [jackson], where simultaneous execution of multiple mechanisms induces equilibrium behavior. However, these works all require full knowledge of value distributions and offer no guarantees under off-equilibrium behavior.
The second stream, initiated by [gorokh2019remarkable], shifts focus to robust, individual-level guarantees without assuming distributional knowledge. They introduce the notion of an agent’s ideal utility and show that a simple threshold strategy in a repeated first-price auction with artificial credits guarantees each agent at least half of their ideal utility. This $1/2$-robustness guarantee has since been matched and extended: [banerjee2023robust] provide a simpler reserve-price mechanism, and [fikioris2023online] match the guarantee via Dynamic Max-Min Fairness (DMMF), a non-market mechanism that allocates to the agent with the least normalized wins. DMMF even achieves stronger guarantees under mild distributional assumptions. However, recent work by [chido] shows that DMMF lacks equilibrium support under natural threshold strategies, even with just two agents, thereby raising questions about its practical relevance. More recently, [lin2025online] improved the $1/2$ factor to $2-\sqrt{2}\approx 0.59$ by randomizing the amount of artificial credits bid when above the threshold (but again, with no equilibrium guarantees).
Our approach draws on two key tools: repeated all-pay auctions [gorokh2021monetary] and the use of Border’s theorem [border1991implementation, border2007reduced] to implement interim allocations. While Border’s theorem has been instrumental in algorithmic mechanism design [cai2012optimal, cai2012algorithmic, alaei2012bayesian, bhalgat2013optimal], our use is existential rather than algorithmic—we strengthen the theorem to establish the existence of robust equilibria. This is closest in spirit to [che2013generalized], who modify allocation mechanisms to enforce lower and upper bounds on allocation probabilities. However, our setting is more complex: we must modify the Border polytope to penalize arbitrary agent behavior while preserving feasibility, a challenge not present in their capacity-based formulation.
@@ 2 Preliminaries
We consider the following simple setting of repeatedly allocating a single, indivisible item among $n$ agents; this was first introduced in the work of [guo2009competitive], and has since been widely studied, as we discuss in subsection˜1.3. At each discrete time-step $t=1,2,\dots,T$, a principal receives a new item and must select an agent $i\in[n]$ to allocate the item to (or not allocate the item).
@@ 2.1 Ideal Utility and Benchmarks
As in previous work in this setting [gorokh2019remarkable, banerjee2023robust, fikioris2023online], we assume each agent $i$ has an exogenously defined fair share $\alpha_{i}$, where $\alpha_{i}\geq 0$ and $\sum_{i=1}^{n}\alpha_{i}=1$. The fair share $\alpha_{i}$ of agent $i$ represents the fraction of rounds we want agent $i$ to win the item under ideal circumstances. As done by previous work, we use the benchmark of ideal utility. The ideal utility of an agent $i$ is the maximum expected utility they can obtain from a single round if they can obtain the item simply by requesting it, but they are only allowed to request it with probability at most their fair share $\alpha_{i}$. Formally, the (per-round) ideal utility is the following.
@@
We seek to define a mechanism $M$ where agent $i$ can achieve high average expected utility $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]]$. Specifically, we want that agent $i$ to achieve a high fraction of her ideal utility, i.e., $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]]\geq\lambda_{i}v_{i}^{\star}$ for some $\lambda_{i}$ as large as possible. As discussed in [gorokh2019remarkable], with additional knowledge of valuation distributions, such guarantees naturally translate into efficiency guarantees (by setting fair shares to maximize the ex-ante welfare relaxation). However these guarantees are equally compelling as individual-level guarantees in non-monetary settings, characterizing what fraction of her top rounds each agent $i$ can realize while leaving the resource on at least $(1-\alpha_{i})T$ days for the others to use.
We study what fraction of her ideal utility an agent $i$ can guarantee in two settings: First, we examine the performance of Bayes-Nash equilibria, where each agent is trying to maximize their own utility. Second, we examine what fraction of her ideal utility agent $i$ can guarantee robustly, i.e., even if the other agents $j\neq i$ are playing adversarially and collude to harm agent $i$ without regard for their own utilities. Critically, we want an agent to achieve both the above with the same (ideally simple) strategy. As in section˜1, we want a mechanism that achieves the following.
$\lambda_{\mathtt{NASH}}$-Nash $\lambda_{\mathtt{ROB}}$-robust equilibria: A mechanism with policy profile $(\pi_{1},\pi_{2},\ldots,\pi_{n})$ such that $(\pi_{1},\pi_{2},\ldots,\pi_{n})$ forms a Nash equilibrium as $T\to\infty$ in which each agent $i$ gets a $\lambda_{\mathtt{NASH}}$ fraction of their ideal utility. Policy $\pi_{i}$ is $\lambda_{\mathtt{ROB}}$-robust with $\lambda_{\mathtt{ROB}}$ not much smaller than $\lambda_{\mathtt{NASH}}$: it guarantees agent $i$ a $\lambda_{\mathtt{ROB}}$ fraction of her ideal utility even if agents $j\neq i$ collude and act adversarially.
To understand how the ideal utility benchmark behaves under Bayes-Nash equilibria, we first present a bound on what fraction $\lambda_{\mathtt{NASH}}$ of ideal utility agents can get under ideal conditions. This will help later certify the minimax optimality of our mechanism in the equilibrium setting. Note that this bound is mechanism-agnostic, as in our example below, the mechanism has access to the agents’ realized values.